OpenStudy (anonymous):

Find the perimeter of a quadrilateral with vertices at L (−4, 0), M (4, 6), N (7, 2), O (−1, −4). Round your answer to the nearest hundredth when necessary. 30 units 42 units 48 units 50 units

OpenStudy (anonymous):

@aaronq @e.mccormick @kirbykirby @Miracrown

OpenStudy (anonymous):

@Iramos1998

OpenStudy (kirbykirby):

You can apply the distance formula for LM, MN, NO, and OL using \(distance = \sqrt{(x_2-x_1)^2+(y_2-y_1)^2}\) Then add up those distances

OpenStudy (anonymous):

I got 250

OpenStudy (anonymous):

@kirbykirby

OpenStudy (kirbykirby):

Oh I ink I see what you did, but actually what you have is: \[ \sqrt{100}+\sqrt{25}+\sqrt{100}+\sqrt{25}=10+5+10+5\]